OpenStudy (anonymous):

How to find the region between y=x and y=2xsqrt(1-x^2) ans: 5/24

OpenStudy (anonymous):

@zepdrix

zepdrix (zepdrix):

lol that teddy bear is awesome c:

OpenStudy (anonymous):

thanks, that's Rosho bear!

zepdrix (zepdrix):

Hmm this is a weird function. Here is what it looks like graphed. https://www.desmos.com/calculator/iqh4upoyle We need the points of intersection. Were you able to find those yet?

OpenStudy (anonymous):

The issue I'm having with this problem is finding the exact intersection of the two lines.

zepdrix (zepdrix):

ah ok :)

zepdrix (zepdrix):

So we set the two functions equal to one another, Then solve for x.\[\large y=x \qquad \qquad \qquad y=2x\sqrt{1-x^2}\] \[\large x=2x\sqrt{1-x^2}\]Solve for x! :D

OpenStudy (anonymous):

right i'm a complete idiot...

zepdrix (zepdrix):

I blame the bear -_-

OpenStudy (anonymous):

Well, the bear is smarter than I am.